The weekly demand for sandwiches at a local sandwich shop is given by:

Qd = 2,000 – 5P + 2Pj – 0.01INCOME,

            where Qd is the number of sandwiches demanded per week, is the per-sandwich price, Pj is the price of a related product, and INCOME is the average monthly income of consumers. 

Suppose that = $10, Pj = $50and INCOME = $5,000. What is the value of the cross-price elasticity of demand? Is the related product a substitute good or a complement?

Suppose that = $10, Pj = $50and INCOME = $5,000. What is the value of the income elasticity of demand? Are sandwiches a normal or an inferior good?



1
Expert's answer
2022-03-03T14:58:56-0500

"Q= 2000- 5(10)+2(50)-0.01(5000)= 2000"

"IE=\frac{ \Delta Q}{\Delta P}\times \frac{I}{Q}"

"-0.01\times \frac{5000}{2000}= -0.025"

This an inferior good

Cross Elasticity"=" "\frac{ \Delta Q}{\Delta P}\times \frac{P}{Q}"

"2\times \frac{50}{2000}= 0.05"

This is a compliment good